How do you write the ration as a fraction in the lowest term 1 3/11 to 1 7/8
Answer by rothauserc(4718)

1 3/11 = 14/11
1 7/8 = 15/8
now we are asked for the ratio
(14/11) / (15/8) = 14/11 * 8/15 = 112 / 165